# ViTally Consistent: Scaling Biological Representation Learning for Cell Microscopy

### v1

@@ Abstract
Large-scale cell microscopy screens are used in drug discovery and molecular biology research to study the effects of millions of chemical and genetic perturbations on cells. To use these images in downstream analysis, we need models that can map each image into a feature space that represents diverse biological phenotypes consistently, in the sense that perturbations with similar biological effects have similar representations. In this work, we present the largest foundation model for cell microscopy data to date, a new 1.9 billion-parameter ViT-G/8 MAE trained on over 8 billion microscopy image crops. Compared to a previous published ViT-L/8 MAE, our new model achieves a 60% improvement in linear separability of genetic perturbations and obtains the best overall performance on whole-genome biological relationship recall and replicate consistency benchmarks. Beyond scaling, we developed two key methods that improve performance: (1) training on a curated and diverse dataset; and, (2) using biologically motivated linear probing tasks to search across each transformer block for the best candidate representation of whole-genome screens. We find that many self-supervised vision transformers, pretrained on either natural or microscopy images, yield significantly more biologically meaningful representations of microscopy images in their intermediate blocks than in their typically used final blocks. More broadly, our approach and results provide insights toward a general strategy for successfully building foundation models for large-scale biological data. [^1] Correspondence: kian.kd@recursion.com, info@rxrx.ai
@@ 1 Introduction
Large-scale cell microscopy assays are used to discover previously unknown biological processes (Przybyla & Gilbert, 2022; Bock et al., 2022; Rood et al., 2024) and identify novel drug candidates and targets (Vincent et al., 2022). Labs are now able to achieve extremely high throughput by leveraging high content screening (HCS) systems that combine automated microscopy with robotic liquid handling (Boutros et al., 2015). Extracting meaningful features from microscopy images in large-scale screens has become increasingly difficult as this scale has increased. Public datasets like RxRx3 (Fay et al., 2023) and JUMP-CP (Chandrasekaran et al., 2023) now include millions of cellular images across 100,000s of unique chemical and genetic perturbations. In addition to limitations in expressiveness of the features that can be derived from them, traditional methods relying on customized pipelines for segmentation, feature extraction, and downstream analysis (Caicedo et al., 2017) struggle to handle this scale effectively  (Chandrasekaran et al., 2021; Carpenter et al., 2006).
The size and complexity of large-scale microscopy data demands image models that can extract rich biological features and do so consistently across experimental replicates, both of which are crucial for downstream biomedical applications. Rich, biologically meaningful representations reveal relationships between genes or compounds to drive the discovery of novel targets and drug candidates, while consistency in features extracted across replicates ensures that findings are reproducible and reliable for therapeutic development.
Foundation models have been developed for representing high-dimensional unstructured biological data such as protein structures (Jumper et al., 2021) and transcriptomics (Hao et al., 2024), but the scale and dimensionality of large-scale microscopy data present unique challenges for generating representations that are both biologically informative and consistent across replicates. HCS datasets are often confounded by complex noise known as batch effects (Caicedo et al., 2017), stemming from differences between experimental batches and biological variability. These batch effects – including natural variation in cell populations – obscure the biological effects of perturbations and make it challenging to isolate the specific effects of the perturbations applied (Yang et al., 2019). Overcoming these obstacles with a model capable of generating robust, biologically meaningful representations can empower HCS to systematically interrogate gene function and identify novel drug candidates (Rood et al., 2024).
State-of-the-art (SOTA) deep learning methods for microscopy leverage Vision Transformers (ViT) (Dosovitskiy et al., 2020) trained with self-supervised learning (SSL) techniques (Balestriero et al., 2023) to learn unbiased representations from large-scale screens (Doron et al., 2023; Kim et al., 2023; Bourriez et al., 2024). Recent studies have demonstrated that ViTs trained as Masked Autoencoders (MAEs) (He et al., 2022) can effectively scale beyond previous approaches and outperform various supervised and smaller SSL models in capturing biologically informative representations of cell images (Kraus et al., 2024). However, the level of consistency found in these representations across a large number of experimental replicates was not previously reported. Furthermore, compared to recent multi-billion parameter transformers developed for natural images (Dehghani et al., 2023) and natural language (Llama3, 2024), model scale in microscopy lags behind (Kraus et al., 2024; Chen et al., 2023a) despite the existence of massive datasets.
In this work, we developed the largest foundation model to date for cell microscopy images, achieving SOTA results in both replicate consistency and biological recall of known gene-gene relationships. Specifically our work offers the following contributions:
  • We demonstrate that training on a curated microscopy dataset of statistically significant positive samples, named Phenoprints-16M, improves both recall of known gene-gene relationships and consistency of embeddings for gene knockout perturbations (Figure 1A). We describe components of this curation strategy that can be generalized to other scientific datasets (§ 3.1).
  • We present a new foundation model, MAE-G/8, a 1.86 billion parameter ViT-G/8 MAE trained on Phenoprints-16M over 48,000 H100 GPU hours on more than 8 billion samples from the curated dataset (Figure 1A, § 3.2).
  • We propose new set of biological linear probing tasks to evaluate representations learned by intermediate ViTs blocks for microscopy data (§ 4). Performance on these linear probing tasks are strongly correlated with performance on important whole-genome scale evaluation metrics while requiring significantly less resources to compute (Figure 4).
  • We find that using intermediate layers leads to better performance on these downstream whole-genome benchmarks at a lower computational inference cost, across SSL ViTs trained on microscopy or natural images. By taking advantage of our linear probing proxy task, we are able to cheaply find the best performing intermediate block (Eq. 1).
Our results indicate that the biological scaling properties first identified by Kraus et al. (2023) extend to the multi-billion parameter regime (§ A.9). We show that our MAE-G/8 model produces a nearly 60% more phenotypically linearly separable latent space compared to previous approaches using the final block of MAE-L/8 (Figure 4), correlating with significant improvements in both recall and replicate consistency when benchmarking across the whole genome (Figure 1B).
@@
Figure 1: (A) Overview of performance gain from different foundation model pretraining and inference strategies. (B) Example whole-genome results for replicate consistency and biological relationship recall on StringDB for models trained with different combinations of strategies, by model name and dataset (left to right): MAE-L/8 (RPI-93M, block $b=24$), MAE-L/8 trimmed to block $b^{*}=15$, MAE-L/8 (Phenoprints-16M, block $b^{*}=20$), MAE-G/8 (Phenoprints-16M, $b^{*}=38$), where $b^{*}$ is the optimal block according to linear probes as defined in Equation 1.
@@ 2 Related work
Recent work suggests that intermediate layers (or, blocks) in large ViTs may achieve superior performance on certain linear probing tasks compared to the final encoder layer (Evci et al., 2022; Dehghani et al., 2023). For example, Alkin et al. (2024) reported that intermediate layers in large MAE-ViTs (ViT-L, ViT-H) have superior ImageNet-1K k-NN accuracy. They attributed this property to the later encoder layers becoming more optimized for the reconstruction task.
@@ 3.1 Training Dataset Curation
To address this, we constructed an aggressively curated training dataset (§ A.1). To learn an initial representation, we began by reproducing the MAE-L/8 model of Kraus et al. (2024) on a dataset of similar size consisting of 93 million HCS images. Using this representation, we first filtered perturbations that did not induce consistent morphological changes to cells. To perform this filtering, we utilized Celik et al. (2024)’s non-parametric perturbation consistency test (§ A.3) Typical Variation Normalization (Ando et al., 2017a; Kraus et al., 2024). This test was applied within each experiment for computational efficiency, and we restricted the analysis to wells containing single perturbations. This consistency was computed for CRISPR guides, siRNAs, and particular concentrations of small molecules across replicates of the same perturbation. P-values were computed for each gene and each (perturbation, concentration) pair. When multiple experiments existed for the same condition, we combined p-values using the Cauchy Combination test (Liu & Xie, 2018).
@@ 3.2 Models
Model Name Parameters Blocks Model Dim Pretraining Data
Baselines
Untrained ViT-S/16 25M 12 384 N/A
Dino-V2 ViT-S/14 25M 12 384 Natural images
Dino-V2 ViT-L/14 307M 24 1024 Natural images
Dino-V2 ViT-G/14 1,100M 40 1536 Natural images
ViT-L/16 WSL 307M 24 1024 Imagenet-21k
ViT-L/16 MAE 307M 24 1024 Imagenet-21k
MAEs for microscopy
CA-MAE-S/16 25M 12 384 RxRx3
MAE-L/8 307M 24 1024 RPI-93M
MAE-L/8 307M 24 1024 Phenoprints-16M
MAE-G/8 1,860M 48 1664 Phenoprints-16M
@@
We compare to several non-finetuned baseline ViT image encoders: three different Dino-v2 backbones (Oquab et al., 2024) (with 4 register tokens (Darcet et al., 2024)) trained on a curated non-biological natural image dataset; a weakly supervised (WSL) classifier ViT-L/16 trained on Imagenet-21k (Ridnik et al., 2021); a MAE ViT-L/16 trained on Imagenet-21k (He et al., 2022); and an untrained ViT-S/16. Preliminary investigations found that channel-wise self-standardization worked best as the image normalization preprocessing for these baselines, and that the class token was slightly better than the global pool of the patch tokens (except for MAE). Convolutional weights in the patch embedding layer were repeated to embed 6 channel images when using models trained on RGB datasets (Wightman, 2019).
@@
Our primary point of comparison is with respect to the best pretrained foundation model presented by Kraus et al. (2024), the MAE-ViT-L/8+ trained on RPI-93Mrained for approximately 40 epochs, learning from over 3.5 billion image crops, using the L2 mean squared error loss function plus an additional Fourier domain reconstruction loss term.
@@
Holding the model backbone constant compared to the MAE-ViT-L/8 by Kraus et al. (2024), we assess the impact of our curated dataset in contrast to the 93M dataset by training a new ViT-L/8 MAE for 500 epochs on Phenoprints-16M.
@@
Figure 3: Block-wise validation set linear probe results comparing ViT models pretrained on cell microscopy images (left) versus natural images (right). (a) 1139-class RxRx1 SiRNA knockdown classification (Sypetkowski et al., 2023); (b) 40-class Anax functional gene group classification on HUVEC cell images from RxRx3 CRISPR knockouts (Fay et al., 2023).
@@ 4 Linear probing representation learning across ViT blocks
Figure 4: Correlations between validation set linear probing (Figure 3) on Anax and RxRx1 for best and last blocks (Eq. 1) compared to downstream whole-genome benchmarks (Table 2) for biological relationship recall on StringDB at 0.05-0.95 threshold and replicate consistency KS statistic.
@@
We expect high quality representations of cell images to generate similar embeddings for cells with the same perturbation, hence a simple linear probe should be able to predict gene perturbation from these representation reasonably well. We train linear probes on the publicly-available RxRx1 dataset in Sypetkowski et al. (2023) which consists of 125,510 high-resolution fluorescence microscopy images of human cells under 1,138 siRNA-induced gene knockdowns (plus unperturbed controls) across four cell types (HEPG2, HUVEC, U2OS, RPE). These gene knockdowns produce strong phenotypes which makes the prediction task more feasible.
@@
Biologically meaningful representation of microscopy images of genetically perturbed cells should capture functional relationships between genes, hence a simple linear probe should be able to predict functional gene groups when trained on these representations. We curated a small subset of 80,000 wells from RxRx3 (Fay et al., 2023) to evaluate linear probes on functional group prediction. We also evaluated similar whole genome knockout screens with ARPE-19 and an additional population of HUVEC cells with soluble TNF-$\alpha$ added to all wells. We manually curated Anax, a set of 40 functionally-diverse gene groups containing 348 genes, with details provided in (§ A.8). Examples of groups include major protein complexes (e.g. proteasome, ribosome-small/large), metabolic pathways (e.g. Krebs cycle) and signaling pathways (e.g. calcium signaling) (Figure 2). These groups span broad biological processes that are conserved across cell types – linear separability of these groups would likely indicate that representations are biologically meaningful regardless of cell type.
As shown in 3(b), MAE-G/8 significantly outperforms other models in Anax group linear probe classification. The best representations once again are obtained from an intermediate block, achieving a balanced accuracy ($0.32$) that is $5\%$ greater compared to its final block’s output features. We observed similar trends for ViT models pretrained on natural images and representations computed from microscopy images of other cell types/conditions (§ A.5, Figure 6).
@@ 5 Whole-genome benchmarking
Table 2: Multivariate known biological relationship recall and univariate replicate consistency benchmarks by model, encoding block $b$, benchmark database, and test statistic. The trimmed models used linear probes to select an earlier block as the feature encoder (Fig. 3). Results are computed over all whole-genome CRISPR knockout perturbation images in RxRx3, after applying TVN and chromosome arm bias correction. For relationship recall, we report results over four databases (React stands for Reactome-PPI (Gillespie et al., 2021)). Best overall result is in bold.
Model backbone $b$ CORUM hu.MAP React StringDB KS CVM
@@
ViT-S/16, Untrained $12$45 .34 .205 .36 .30 4.3
ViT-S/14, Dino-V2 $12$48 .345 .20 .38 .34 5.6
trimmed $5$51 .36 .21 .40 .35 6.0
ViT-L/16, ImageNet WSL $24$52 .35 .21 .39 .34 5.5
ViT-L/14, Dino-V2 $24$49 .34 .21 .38 .34 5.3
trimmed $12$55 .37 .22 .41 .36 5.9
ViT-L/16, ImageNet MAE $24$53 .355 .215 .40 .34 5.1
trimmed $11$53 .36 .22 .40 .35 5.8
ViT-G/14, Dino-V2 $40$44 .31 .20 .35 .29 3.8
trimmed $16$53 .35 .22 .40 .33 5.2
MAEs for microscopy
CA-MAE-S/16 , RxRx3 $12$55 .37 .23 .43 .47 10.4
MAE-L/8 , RPI-93M $24$61 .43 .25 .47 .52 12.3
trimmed $15$60 .43 .255 .475 .57 15.2
MAE-L/8 , PP-16M $24$60 .43 .255 .48 .59 16.2
trimmed $20$60 .435 .26 .48 .59 16.2
MAE-G/8 , PP-16M $48$62 .44 .26 .49 .60 16.4
trimmed $38$615 .44 .26 .49 .63 18.2
@@
Table 2 presents our benchmarks computed across the whole-genome. These evaluate the genomic representations obtained for each model by aggregating millions of embeddings of cell images spanning $>$100,000 of genetic knockout perturbations (17,063 genes $\times$ 6 single guide RNAs each) on HUVEC cells from RxRx3 (Fay et al., 2023). Computing these benchmarks for HCS screens typically requires inferring 140 million crops from the genome-wide RxRx3 microscopy screen (Kraus et al., 2023) (64 tiled crops per each of the 2.2 million wells), but, to reduce compute costs, we discard the outer ring of crops, leaving the 36 center non-edge crops for each well. This requires 80 million forward passes to comprehensively evaluate a new encoder. After inference, we use typical variation normalization (Ando et al., 2017b) and chromosome arm bias correction (Lazar et al., 2023) to post-process the embeddings and aggregate them to the gene-level.
@@
We present the multivariate biological relationship recall benchmarks proposed by Celik et al. (2024) and originally evaluated for MAEs by Kraus et al. (2023, 2024). These metrics evaluate how many annotated pair-wise relationships are recalled from public databases (CORUM, hu.MAP, Reactome-PPI, StringDB) in the extremities of a ranked list of cosine similarities of all pair-wise post-processed embeddings (details in § A.6). To ensure embeddings represent technical replicates of perturbations consistently, we also evaluate model performance on replicate consistency based on the experimental design used in the RxRx3 dataset. Specifically, we compare the similarity of the embedding for corresponding wells across different experiments via a non-parametric statistical test. The test statistic measures the difference between the perturbation replicates’ similarity distribution and an empirical null distribution, with larger values indicating greater consistency (details in § A.7).
In order to compare models, we summarize the resulting statistics over all technical replicates in RxRx3 by taking their median, as reported in columns KS and CVM in Table 2, and visualized in Figure 5. MAEs pretrained on microscopy data show improved performance compared the baseline models. Furthermore, training on the Phenoprints-16M dataset improves the performance of the MAEs significantly and trimmed MAE-G/8 achieves the best overall performance.
Our linear probing analysis (Figure 3) allowed us to trim our models to better encoding blocks. Comparing models on their best respective blocks, MAE-G/8 improves on MAE-L/8 with a 16% improvement in Anax functional gene group classification (.27$\rightarrow$.31) and a 24% improvement in RxRx1 perturbation classification (.41$\rightarrow$.51). Compared to the best published result for whole-genome benchmarks (MAE-L/8 trained on RPI-93M (Kraus et al., 2023)), MAE-G/8 obtains a 20% improvement in replicate consistency KS (.52$\rightarrow$.63) and 4.3% improvement in StringDB recall (.472$\rightarrow$.492). When using our linear probes to select outputs from block $b^{*}=15$ (Equation 1) from that MAE-L/8, the gain for MAE-G/8 changes to 9.2% in KS and 3.5% in StringDB recall.

### v2

@@ Abstract
Deriving insights from experimentally generated datasets requires methods that can account for random and systematic measurement errors and remove them in order to accurately represent the underlying effects of the conditions being tested. Here we present a framework for pretraining on large-scale microscopy datasets that includes three steps: (1) curating a set of diverse and self-consistent training samples, (2) scaling training of an appropriate foundation model architecture on this dataset, (3) evaluating intermediate layers of the trained model to identify the best representation for downstream tasks. Using this strategy, we present the largest foundation model for cell microscopy data to our knowledge, a new 1.9 billion-parameter ViT-G/8 MAE trained on over 8 billion microscopy image crops. Compared to a previous published ViT-L/8 MAE, our new model achieves a 60% improvement in linear separability of genetic perturbations and obtains the best overall performance on whole-genome relationship recall, batch correction replicate consistency, and compound-gene activity prediction benchmarks. [^1] An earlier version of this work appeared at the NeurIPS 2024 Foundation Models for Science Workshop (Kenyon-Dean et al., 2024).
Machine Learning, ICML, MAE, drug discovery, microscopy, SSL, linear probing, biology, high-content screening, foundation models
@@ 1 Introduction
Microscopy images capture detailed information about a cell’s responses to perturbations: indeed many early biological discoveries were made by biologists looking through microscopes studying how cells behave  (Hooke, 1665). But with the advent of high throughput screening platforms that capture cells’ responses to tens or hundreds of thousands of perturbations, we can no longer rely on people to study these images. Instead, we need representations of the cell images that capture the biologically meaningful information and capture the similarities and differences in cells’ response to perturbations.
@@
Figure 1: (A) Overview of performance gain from different MAE pretraining and inference strategies. (B) Example whole-genome results for replicate consistency and biological relationship recall on StringDB for models trained with different combinations of strategies, by model name and dataset (left to right):
State-of-the-art (SOTA) deep learning methods for learning representations of microscopy leverage Vision Transformers (ViT) (Dosovitskiy et al., 2020) trained with self-supervised learning (SSL) techniques (Balestriero et al., 2023) from large-scale screens (Doron et al., 2023; Kim et al., 2023; Bourriez et al., 2024). They have been shown to be very effective at representing the subtle changes in cell morphology in response to perturbations, and masked autoencoders (He et al., 2022) in particular exhibit scaling laws for recovering known biological relationships (Kraus et al., 2024). However, given that this data has to be collected experimentally, there is a limit to how far we can scale these approaches before one runs out of data. To continue to improve these representations, we need to use experimental data as efficiently as possible. In this paper, we show how curated training datasets combined with careful probing of model representations allows us to scale to larger models that significantly improve recall and separation of biological phenotypes.
Careful curation of data has enabled better data efficiency in language models (Llama3, 2024; Eldan & Li, 2023; Javaheripi et al., 2023). However, unlike language, we cannot simply rely on prior knowledge about the data source to weigh data (e.g. by up-weighting trusted sources like arXiv or wikipedia). Instead, we curate a training dataset that is relatively uniform with respect to the observed phenotypes. Because many perturbations will not induce a morphological change, a random sample of images across all perturbations will be dominated by cells that look like unperturbed cells. We show that one can use weaker models to filter images to perturbations that induce a distinct morphological phenotype, thereby reducing the absolute training set size but increasing the relative diversity. The resulting dataset, Phenoprints-16M, has 5×5$\times$5 × fewer images than the dataset used to train Kraus et al.’s masked autoencoder, but we find that it leads to better performance when combined with scaled compute through longer self-supervised training.
Data curation gives one way of improving a model, but we can also optimize over the layer that we use to construct a representation. Traditionally, the finally hidden layer has been used (Doron et al., 2023; Kim et al., 2023; Kraus et al., 2024), but there is some evidence from mechanistic interpretability (Templeton et al., 2024) that earlier layers may learn more abstract features, which may provide more useful features for biological tasks. We show that using a linear probing proxy task, we are able to cheaply find the best performing intermediate block for many models. We find that using intermediate layers leads to better performance on downstream whole-genome relationship recall and perturbational replicate consistency at a lower computational inference cost, and that these results are consistent across a variety SSL ViTs trained on either microscopy or natural images. Our proxy task involves fitting a set of biological linear probing tasks to evaluate representations learned by intermediate ViTs blocks for microscopy data (§ 4). We find this task correlates extremely well with biological recall benchmarks which allows us to reliably probe the model’s representations.
In summary, we present a series of data curation and model probing techniques that allow us to more effectively use experimental microscopy data. By combining these insights, we are able to train a new foundation model, MAE-G/8, a 1.9 billion parameter ViT-G/8 MAE trained on Phenoprints-16M for 48,000 H100 GPU hours on more than 8 billion samples drawn from the curated dataset (Figure 1A, § 3.2) resulting in significant improvements across a range of challenging biological benchmarks. These results indicate that the scaling properties first identified by Kraus et al. (2023) extend to the multi-billion parameter model regime across a wide variety of newly examined biologically-motivated benchmark tasks.
@@ 2 Related work
Recent work suggests that intermediate layers (or, blocks) in large ViTs may achieve superior performance on certain linear probing tasks compared to the final encoder layer (Evci et al., 2022; Dehghani et al., 2023). Alkin et al. (2024) reported that intermediate layers in large MAE-ViTs (ViT-L, ViT-H) have superior ImageNet-1K k-NN accuracy, likely because later encoder layers become more optimized for the reconstruction task.
Evaluating Representations for Drug Discovery.
@@ 3.1 Training Dataset Curation
To address this, we constructed an aggressively curated training dataset (§ A.1). To learn an initial representation, we began by reproducing the MAE-L/8 model of Kraus et al. (2024) on a dataset of similar size consisting of 93 million HCS images. Using this representation, we first filtered perturbations that did not induce consistent morphological changes to cells. To perform this filtering, we utilized Celik et al. (2024)’s non-parametric perturbation consistency test (§ A.3) after correcting for batch effects using Typical Variation Normalization (Ando et al., 2017; Kraus et al., 2024). This test was applied within each experiment for computational efficiency, and we restricted the analysis to wells containing single perturbations. This consistency was computed for CRISPR guides, siRNAs, and particular concentrations of small molecules across replicates of the same perturbation. P-values were computed for each gene and each (perturbation, concentration) pair. When multiple experiments existed for the same condition, we combined p-values using the Cauchy Combination test (Liu & Xie, 2018).
@@ 3.2 Models
We compare to several non-finetuned baseline ViT image encoders: three different Dino-v2 backbones (Oquab et al., 2024) (with 4 register tokens (Darcet et al., 2024)) trained on a curated non-biological natural image dataset; a MAE ViT-L/16 trained on Imagenet-21k (He et al., 2022); and an untrained ViT-S/16. We found that channel-wise self-standardization worked best as the image normalization preprocessing for these baselines, and that the class token was slightly better than the global pool of the patch tokens (except for MAE). Convolutional weights in the patch embedding layer were repeated to embed 6 channel images when using models trained on RGB datasets (Wightman, 2019).
@@
Our primary point of comparison is with respect to the best pretrained foundation model presented by Kraus et al. (2024), the MAE-ViT-L/8+ trained on RPI-93M. This MAE-L/8 was trained for approximately 40 epochs, learning from over 3.5 billion image crops, using the L2 mean squared error loss function plus an additional Fourier domain reconstruction loss term.
@@
Holding the model backbone constant compared to the MAE-ViT-L/8 by (Kraus et al., 2024), we assess the impact of our curated dataset in contrast to the 93M dataset by training a new ViT-L/8 MAE for 500 epochs on Phenoprints-16M.
@@
Figure 3: Block-wise validation set linear probe results comparing ViT models pretrained on cell microscopy images (left) versus natural images (right); note the difference in y-axes. (a) 1139-class RxRx1 SiRNA knockdown classification (Sypetkowski et al., 2023); (b) 40-class Anax functional gene group classification on HUVEC cell images from RxRx3 CRISPR knockouts (Fay et al., 2023).
@@ 4 Linear probing representation learning across ViT blocks
Figure 4: Correlations between validation set linear probing (Figure 3) on Anax and RxRx1 for best and last blocks (Eq. 1) compared to downstream whole-genome benchmarks (Table 1) for biological relationship recall on StringDB at 0.05-0.95 gene-gene cosine similarity threshold and replicate consistency KS statistic. Models with bold borders are trimmed, red are natural image baseline models and blue are trained on microscopy.
@@
We expect high quality representations of cell images to generate similar embeddings for cells with the same perturbation, hence a simple linear probe should be able to predict gene perturbation from these representation reasonably well. We train linear probes on the publicly-available RxRx1 dataset (Sypetkowski et al., 2023) which consists of 125,510 high-resolution fluorescence microscopy images of human cells under 1,138 siRNA-induced gene knockdowns (plus unperturbed controls) across four cell types (HepG2, HUVEC, U2OS, RPE). These gene knockdowns produce strong phenotypes which makes the prediction task more feasible.
@@
[…]
As shown in 3(b), MAE-G/8 significantly outperforms other models in Anax group linear probe classification. The best representations once again are obtained from an intermediate block, achieving a balanced accuracy ($0.32$) that is $5\%$ greater compared to its final block. We observed similar trends for ViT models pretrained on natural images and representations computed from microscopy images of other cell types/conditions (§ A.5, Figure 6).
@@ 5 Whole-genome benchmarking
Table 1: Multivariate known biological relationship recall and univariate replicate consistency benchmarks by model, encoding block $b$, aggregated recall over a total of 145,447 possible gene-gene relationships annotations accumulated across five benchmark databases, and replicate consistency test statistics. The trimmed models used linear probes to select an earlier block as the feature encoder (Fig. 3). Each benchmark is computed over TVN-aligned gene-aggregated model embeddings, for each one higher is better, and the best overall result is in bold. We report the Recall % of biological relationships obtained below the 5th and above the 95th percentiles (0.05-0.95) of all gene-gene cosine similarities against the null distribution, as per (Celik et al., 2024; Kraus et al., 2024), with mean $\pm$ the standard deviation computed over 3 different random seeds of sampling the null distribution in each benchmark run. For replicate consistency, we report the Komologorov-Smirnov (KS) and Cramer-von Mises (CM) test statistics.
[…]
@@
ViT-S/16 Untrained $12$ 34.2 $\pm$ .08 .30 4.3
ViT-L/16 ImgNet MAE $24$ 37.4 $\pm$ .06 .34 5.1
trimmed $11$ 37.8 $\pm$ .05 .35 5.8
Baseline Dino ViTs
ViT-S/14 Dino-V2 $12$ 35.6 $\pm$ .09 .34 5.6
trimmed $5$ 37.2 $\pm$ .16 .35 6.0
ViT-L/14 Dino-V2 $24$ 35.7 $\pm$ .05 .34 5.3
trimmed $12$ 38.8 $\pm$ .04 .36 5.9
ViT-G/14 Dino-V2 $40$ 32.7 $\pm$ .10 .29 3.8
trimmed $16$ 37.5 $\pm$ .13 .33 5.2
Microscopy MAEs
CA-MAE-S/16 RxRx3 $12$ 39.6 $\pm$ .10 .47 10.4
MAE-L/8 RPI-93M $24$ 44.4 $\pm$ .12 .52 12.3
trimmed $15$ 44.3 $\pm$ .12 .57 15.2
MAE-L/8 PP-16M $24$ 44.4 $\pm$ .12 .59 16.2
trimmed $20$ 44.7 $\pm$ .06 .59 16.2
MAE-G/8 PP-16M $48$ 45.4 $\pm$ .07 .60 16.4
trimmed $38$ 45.4 $\pm$ .15 .63 18.2
@@
Table 1 presents our benchmarks computed across the whole-genome. These evaluate the genomic representations obtained for each model by aggregating millions of embeddings of cell images spanning $>$100,000 of genetic knockout perturbations (17,063 genes $\times$ 6 single guide RNAs each) on HUVEC cells from RxRx3 (Fay et al., 2023). Computing these benchmarks for HCS screens typically requires inferring 140 million crops from the genome-wide RxRx3 microscopy screen (Kraus et al., 2023) (64 tiled crops per each of the 2.2 million wells), but, to reduce compute costs, we discard the outer ring of crops, leaving the 36 center non-edge crops for each well. This requires 80 million forward passes to comprehensively evaluate a new encoder. After inference, we use typical variation normalization (Ando et al., 2017) and chromosome arm bias correction (Lazar et al., 2024) to post-process the embeddings and aggregate them to the gene-level.
@@
We present the multivariate biological relationship recall benchmarks proposed by Celik et al. (2024) and originally evaluated for MAEs by Kraus et al. (2023, 2024). These metrics evaluate how many annotated pair-wise relationships are recalled from public databases (CORUM, hu.MAP, Reactome-PPI, Signor, StringDB) in the extremities of a ranked list of cosine similarities of all pair-wise post-processed embeddings (details in § A.6).
To ensure embeddings represent technical replicates of perturbations consistently, we also evaluate model performance on replicate consistency based on the experimental design used in the RxRx3 dataset. Specifically, we compare the similarity of the embedding for corresponding wells across different experiments via a non-parametric statistical test. The test statistic measures the difference between the perturbation replicates’ similarity distribution and an empirical null distribution, with larger values indicating greater consistency (§ A.7). To compare models, we summarize the resulting statistics over all technical replicates in RxRx3 by taking their median, reported in columns KS and CM (Kolmogorov–Smirnov and Cramer-von Mises test statistics) in Table 1.
Not suprisingly, even the smallest CA-MAE-S/16 trained on microscopy data outperforms the largest baselines ViTs trained on natural images. Training self-supervised on our novel carefully curated Phenoprints-16M dataset improves the performance of the MAEs, as does trimming to an early layer detected by linear probes, while the most scaled model MAE-G/8 achieves the best overall performance in all respects when trimmed. Compared to the best previously SOTA model at its normally used final block, MAE-L/8 RPI-93M (Kraus et al., 2023), MAE-G/8 trimmed obtains: (a) the highest statistically significant gain in relationship recall over the previous SOTA, achieving a z-score of improvement of 5.21; (b) a 21% improvement in the KS statistic (.52$\rightarrow$.63); and, (c) a 48% improvement in the replicate consistency CM statistic (12.3$\rightarrow$18.2).
Table 2: Comparing MAEs to manually extracted CellProfiler (McQuin et al., 2018) features on biological relationship recall. Reported at 0.05-0.95 cosine threshold on the public JUMP-CP image dataset (Chandrasekaran et al., 2023), which is generated by completely different labs and assay protocols compared to the images used for pretraining. Each result has a standard deviation ±.$\leq\pm.0023$, spanning gene-gene relationships across nearly 8,000 gene-knockouts.
Model backbone $b$ Pretraining data CORUM hu.MAP Reactome StringDB
CellProfiler - N/A .219 .184 .131 .191
CA-MAE-S/16 12 RxRx3 .233 .199 .154 .214
MAE-L/8 24 RPI-93M .248 .208 .160 .226
MAE-G/8 trimmed 38 Phenoprints-16M .264 .215 .165 .235